OpenStudy (anonymous):

Solve for B: A = 3(B + C)

OpenStudy (anonymous):

divide both sides by 3, then youre left with a/3=b+c and then subract c on both sides, and the final answer is (a/3) -c=b :)

OpenStudy (anonymous):

A=3B + 3C 3B=-3C+A B=-3C/3 + A/3